OpenStudy (anonymous):

the pyramids at Giza include several square-based pyramids that were built in honor of the Egyptian gods. One of these incredible monuments is the Khafre pyramid, built as part of a temple complex, with the famous Sphinx at its base. The base length of this pyramid is 705 feet, and its height is 471 feet. Massive dimensions indeed! Calculate the volume of the Khafre pyramid. Show the steps of your solution and explain your work.

OpenStudy (anonymous):

is this correct Volume = (1/3) Bh where B is the area of the base and h is the height V = 1/3 (705)^2 (471) V = 78, 032, 925 cubic feet

OpenStudy (anonymous):

V = (1/3) * h * l^2 V = (1/3) * 471 * 705^2 V = 78,032,925

OpenStudy (anonymous):

volume of a pyramid is (1/3)Ah where A is the area of the ground and h is the height taken from the peak. but it says base length not base surface so ...............

OpenStudy (anonymous):

you would need another side to calculate the area of the base right? and then you can use that in the formula as done above

OpenStudy (anonymous):

thx for the help i think i got it, 91827364 and everyone 2

OpenStudy (phi):

918, it mentions square-based pyramids

OpenStudy (anonymous):

thx @phi i missed that detail